### v1

@@ 4 Study Finding
According to our experimental results in Table 9, existing techniques perform higher in localizing bugs from deep learning libraries than that of frameworks and tools. This could be attributed to the modular design of libraries, aimed at handling specific tasks independently [57]. In the example bug report (Table 10 [60]), the autograd module of the Torch.Library [^7] https://pytorch.org/docs/stable/library.html plays a crucial role in calculating gradient across all tensor operations, whereas autograd.gradcheck [^8] https://pytorch.org/docs/stable/generated/torch.autograd.gradcheck.html is a utility function for verifying the accuracy of these computed gradients. Here, a specific error concerning gradient computation in gradcheck was reported during sparse matrix multiplication. Such problems can be localized more efficiently when occurring within a clearly defined module like autograd.gradcheck. In this context, DNNLOC identified the buggy file at the Top@1 level, while BLUiR located the correct file at Top@3. BLUiR’s performance suggests that libraries’ modular architecture could facilitate the bug localization process, which leverages structural analysis. On the other hand, DNNLOC’s success in accurately locating bugs can be linked to its ability to capture complex patterns using deep learning, which works effectively with the modular design of library bugs.
@@
In contrast, we notice from Fig. 3, the performance of bug localization techniques dropped for framework bugs, especially when using the BLUiR method. Frameworks consist of a broad architecture that guides the flow of control, involving multiple layers and components that interact in complex ways [55, 10]. The example bug in Table 11 [61] deals with CTC Loss in a Keras model with LSTM on a TensorFlow Processing Unit (TPU) highlights a framework-level issue. For this example bug, DNNLOC retrieved the correct buggy file at the top 38 position. Despite DNNLOC’s strength in handling complex relationships, its lower rank demonstrates a shortcoming, suggesting the need to address the complexities within deep learning frameworks better.
BugLocator retrieved the correct buggy file at the top 65 position, indicating its limitations in transcending textual similarities between framework code and bug reports. Finally, BLUiR emerges as the least effective, ranked the correct file at the top 93 position, highlighting its insufficiency in localizing the framework-level bugs. BLUiR may struggle to match the bug report keywords to the correct code segment since the framework’s vast codebase introduces a high degree of variance, making the bug localization challenging.
@@
Lastly, the performance in bug localization slightly improved for tools, but it was not as good as with libraries. This improvement could be attributed to the fact that tools (e.g., TensorBoard) are more independent and have less dependencies than the entire framework (e.g., Tensorflow). While tools like TensorBoard focus on specialized tasks like visualization, they often integrate with complex frameworks like TensorFlow. The example bug in Table 12 refers to an incorrect rendering of floating point metrics in the TensorBoard HParams dashboard. It can be classified as a tool bug, as TensorBoard is a visualization utility within the TensorFlow ecosystem. In this instance, DNNLOC identified the correct buggy file at the top 27 position. BugLocator located the correct file at the 40th position, which highlights its limitations in localizing tool bugs that warrants more than textual analysis. Meanwhile, BLUiR identified the correct file at the top 52 position, reflecting BLUiR’s struggle to effectively match textual descriptions in bug reports with the specific code segments responsible for visual aspects.
@@
[…]
In this research question, we investigate the characteristics and localization challenges of different types bugs in deep learning systems through manual analysis. First, we employ stratified random sampling to construct a random sample that represents a balanced presence of instances from each bug type [63]. We select 385 bugs from both of the subsets of our datasets (BugGL, Denchmark from Table 3) that have 95% confidence interval and 5% error margin.
We performed our manual analysis using 385 bug reports from the Denchmark dataset. We did two levels of classification from deep learning systems. First, we manually labeled the bugs as deep learning-related (DL) bugs and non-deep learning-related (NDL) bugs from the deep learning systems involving two annotators. Second, we labeled the deep learning-related (DL) bugs into five categories: Model, Training, Tensor, API, and GPU, based on the existing taxonomy of Humbatova et al. [30]. We also analyzed the bug reports, associated developers’ discussions, and bug-fix code changes as a part of the labeling. Two authors of this work labeled the sample dataset separately and achieved a Cohen’s kappa [64] of 0.80, which indicates a substantial agreement between the authors. Our manual analysis above was documented using an Excel sheet, with a total of $\approx$55 hours spent by each author, which is provided in our replication package [46].
@@
Table 13: Experimental result of existing bug localization techniques (BugLocator, BLUiR, BLIA, DNNLOC) of each category of bugs in deep learning systems
Method Model Training Tensor API GPU MAP BugLocator 0.368 0.312 0.235 0.446 0.183 BLUiR 0.293 0.403 0.601 0.258 0.266 BLIA 0.357 0.345 0.448 0.395 0.290 DNNLOC 0.593 0.411 0.619 0.493 0.296 MRR BugLocator 0.532 0.386 0.358 0.378 0.223 BLUiR 0.387 0.427 0.621 0.424 0.311 BLIA 0.472 0.419 0.553 0.447 0.327 DNNLOC 0.585 0.467 0.682 0.548 0.336 MAP= Mean Average Precision MRR=Mean Reciprocal Rank
[…]
Figure 5: Performance of existing bug localization techniques (BugLocator, BLUiR, BLIA, DNNLOC) for each type of bug in deep learning systems
@@
Model bugs: From Table 13, we see that DNNLOC performs the best for model bugs, which could be attributed to its ability to capture complex patterns using DNN and textual relevance from the bug reports and source code using rVSM. Model bugs are often connected to a model’s type, properties, and layers. According to our observation, the texts describing model-related issues in bug reports have significant vocabulary overlap with the implementation of a deep learning model.
@@
Table 14 [39] shows an example bug report that discusses a model bug from the CASL.ai project. Fig. 8 [39] (as shown in Appendix A) shows the code snippet responsible for the bug. The bug in the GPT2Tokenizer lies in the _bpe method, causing faulty tokenization and impacting the functionality of the GPT2 language model, which can be considered a model bug. The incorrect character merging during byte pair encoding leads to faulty tokenization.
BugLocator incorrectly retrieves ‘SentencePieceTokenizer.py’ (Fig. 9) (as shown in Appendix A) as the Top@1 result. It relies on lexical overlap between bug reports and source code. We found four main keywords from the bug report (Table 14) —
GPT2Tokenizer, recover, seq2seq, and model — overlapping with an incorrect file (i.e., SentencePieceTokenizer.py). This lexical similarity could have led to an incorrect localization. Interestingly, the ground truth file was retrieved at the 8th position (Fig. 8) (as shown in Appendix A) by BugLocator. BLIA retrieves the same ground truth file at the 17th position, which is less than ideal.
@@
On the other hand, the BLUiR approach retrieves the ground truth code at a lower position (top 131). Thus, the similarity between bug reports and source code elements might not be sufficient for locating model bugs. In the example above (Table 14), the bug is triggered when tokenizing the following sentence – ‘BART is a seq2seq model.’. BLUiR incorrectly retrieved the source code file with the Seq2Seq class at the Top@1 position (Fig. 10) (as shown in Appendix A). One possible explanation could be that several important keywords from the bug report align with similar (e.g., ‘seq2seq’, ‘encode’, and ‘model’) code elements (e.g., seq2seq class), which are relevant to the false positive bug. Such a misalignment occurs due to BLUiR’s heavy reliance on structural elements (e.g., class, method, variable) and less attention to the semantic aspect of the bug report.
Finally, the DNNLOC approach retrieves the buggy file correctly at the Top@1 position for the model bug (Table 14). Unlike IR-based methods, DNNLOC’s neural networks can capture deeper semantic links that go beyond lexical similarities. For instance, it has a better chance to semantically connect terms such as ‘seq2seq’ and ‘tokenizer’ in the bug report to their corresponding code-level implementations in GPT2Tokenizer despite their textual mismatch. Thus, DNNLoc’s capability to leverage the non-linear relationships might have helped it localize model bugs better in deep learning systems.

### v2

@@ 4.1.2 Comparison among the categories of deep learning systems
bjXnet 0.605 0.550
Figure 5: Performance comparison of existing approaches (BugLocator, BLUiR, BLIA, DNNLOC, bjXnet) for various deep learning project classes
According to our experimental results in Table 14, existing techniques perform higher in localizing bugs from deep learning libraries than that of frameworks and tools. This could be attributed to the modular design of libraries, aimed at handling specific tasks independently [48]. In the example bug report (Table 15 [75]), the autograd module of the Torch.Library [^8] https://pytorch.org/docs/stable/library.html plays a crucial role in calculating gradient across all tensor operations, whereas autograd.gradcheck [^9] https://pytorch.org/docs/stable/generated/torch.autograd.gradcheck.html is a utility function for verifying the accuracy of these computed gradients. Here, a specific error concerning gradient computation in gradcheck was reported during sparse matrix multiplication. Such problems can be localized more efficiently when occurring within a clearly defined module like autograd.gradcheck.
Both DNNLOC and bjXnet retrieved the buggy file at Top@1 by leveraging this modularity. DNNLOC’s success in accurately locating bugs can be linked to its ability to capture complex patterns using deep learning, which works effectively with the modular design of library bugs. On the other hand, bjXnet uses attention mechanisms to focus on critical functions or variables, like autograd.gradcheck, which might lead to more precise localization and demonstrate the strength of attention-based techniques in isolating and targeting key components in modular architectures.
@@
In contrast, we notice from Fig.5 that the performance of bug localization techniques dropped for framework bugs, especially when using the BLUiR method. Frameworks consist of a broad architecture that guides the flow of control, involving multiple layers and components that interact in complex ways [46, 10]. The example bug in Table 15 [76] deals with CTC Loss in a Keras model with LSTM on a TensorFlow Processing Unit (TPU), highlighting a framework-level issue. For this example bug, DNNLOC retrieved the correct buggy file at the top 38 position, while BugLocator retrieved it at the top 65 position, indicating limitations in relying only on textual similarities between framework code and bug reports. BLUiR, emerging as the least effective, ranked the correct file at the top 93 positions, highlighting its insufficiency in localizing framework-level bugs with only structural code analysis. bjXnet retrieved the correct buggy file at the top 28 position. All baseline techniques struggle with framework bugs due to the complex, multi-layered architecture of frameworks, which lacks the clear modular boundaries found in libraries. Even DNNLOC and bjXnet face significant challenges because the complex dependencies and interactions across multiple components in frameworks make it harder to trace and isolate the exact location of faults.
@@
Lastly, the performance in bug localization slightly improved for tools, but it was not as good as with libraries. This improvement could be attributed to the fact that tools (e.g., TensorBoard) are more independent and have fewer dependencies than the entire framework (e.g., TensorFlow). While tools like TensorBoard focus on specialized tasks such as visualization, they often integrate with complex frameworks like TensorFlow. The example bug in Table 17 refers to an incorrect rendering of floating point metrics in the TensorBoard HParams dashboard. It can be classified as a tool bug, as TensorBoard is a visualization utility within the TensorFlow ecosystem. In this instance, bjXnet retrieves the correct buggy file at the top 25 position. DNNLOC identified the correct buggy file at the top 27 position, BugLocator located the correct file at the 40th position, and BLUiR identified the correct file at the top 52 position. These results suggest that while tools are more independent, their connection to complex frameworks still creates challenges. bjXnet and DNNLOC perform slightly better by understanding the structure more deeply, but all techniques find it difficult to handle the interaction between tools and frameworks.
@@
4.1.3 Preliminary study on spectrum-based bug localization
Specturm-based bug localization techniques have been widely used for decades in traditional software systems [55]. However, they might face significant challenges when applied to deep learning systems. Spectrum-based bug localization techniques analyze the execution traces of passing and failing tests to measure the frequency with which each program element is executed in both cases [55, 78]. They rank elements by suspiciousness to prioritize fault-prone locations. For example, Tarantula [55] calculates the suspiciousness score based on the ratio of failed to total executions of an element, highlighting those more frequently associated with failures. However, capturing execution traces against passing and failed tests from deep learning systems could be highly challenging. Besides, not all bugs in deep learning systems cause crashes or generate a stack trace [36]. Instead, some lead to incorrect behavior, known as silent bugs [57]. To better understand the challenges of spectrum-based bug localization techniques, we conducted a case study to assess their applicability in deep learning systems.
Figure 6: Model architecture of the selected deep learning model
Bug selection: We randomly selected a training bug from a classification model for our case study (see Fig.6). Here, the final layer of the model is defined as activation=None (buggy) instead of activation=’softmax’ (correct). The bug causes the final layer to output unnormalized logits. Although the cross-entropy loss is still computed, these raw outputs hinder the expected decrease in loss over multiple epochs, thereby hampering model training.
Ground truth for bug localization: The final layer’s activation function is incorrect, which makes the layer buggy (a.k.a., ground truth). We aim to determine whether spectrum-based localization techniques can correctly locate this buggy layer from the model.
Choice of spectrum-based technique: To perform the preliminary study, we chose a seminal work on spectrum-based bug localization – Tarantula [55]. Tarantula uses the execution profiles of test cases to compute suspiciousness scores for program components (e.g., statements). Several contemporary studies [79] have extended Tarantula’s spectra-based approach by incorporating machine learning and deep learning techniques. However, we wanted to directly assess Tarantula’s effectiveness and limitations in locating faults in deep learning systems without introducing any confounding factors from hybrid methods. Hence, we selected Tarantula for our case study.
Figure 7: Top 10 Neuron-Level Suspiousness Scores by Layers
Table 18: Pass/Fail Results from 20 Training Runs
Run Result Run Result Run Result
1 Failed 8 Passed 15 Passed
2 Failed 9 Passed 16 Failed
3 Passed 10 Passed 17 Failed
4 Passed 11 Passed 18 Passed
5 Failed 12 Failed 19 Failed
6 Failed 13 Passed 20 Failed
7 Failed 14 Failed
Test cases: The training of deep learning models are inherently stochastic due to factors such as random weight initialization and data shuffling [80]. Hence, designing test cases for deep learning models is challenging [81]. However, to simulate the traditional approach for spectrum-based bug localization, we needed to create both failing and passing test cases. Since the final layer of our target model is defined without a softmax activation, its output remains unnormalized. This unnormalized output is processed by the model’s loss function (e.g., categorical cross-entropy), but the loss is calculated using raw logits, which is undesirable. However, in some training runs, favorable random initialization and data ordering may cause the loss to decrease, even when the bug (i.e., incorrect activation function) is present. In other cases, the loss might only decrease. Building on prior research [81, 82, 80], we conducted 20 separate training runs, each with 10 epochs and different random seeds, to capture this variability. We define our passing and failing test cases as follows:
- Failing test cases: Training runs where the loss stagnates across $N$ epochs (here, $N=3$, with a total of 10 epochs).
- Passing test cases: Training runs where the loss decreases despite the model containing the fault.
Coverage Metric: Our study captures test coverage at two levels: layer level and neuron level. First, we record whether each layer in the model is executed during each training run. This binary coverage metric aligns with traditional spectrum-based techniques, where suspiciousness scores are calculated based on how components are covered by the passing and failing test cases. Similarly, we calculate the suspiciousness score for a given layer $l$ as follows:
- Failed($l$): Number of failing runs that execute layer $l$.
- Passed($l$): Number of passing runs that execute layer $l$.
Tarantula integrates these statistics into its standard formula to compute a suspiciousness score, ideally flagging layers with a higher ratio of failing-to-passing executions as more suspicious. Second, we record the activation of individual neurons within each layer to assess the contribution of neurons to the overall model behavior. Here, a neuron is considered covered if it is activated (i.e., its output is greater than zero) during a training run. We calculate a Tarantula score for each neuron based on its activation frequency in failing versus passing test cases. Our approach is conceptually similar to the granular coverage criteria introduced by Ma et al. [81].
Table 19: Suspiciousness Scores Per Layer
Layer Name Suspiciousness Score
conv2d 0.5000
conv2d_1 0.5000
max_pooling2d 0.5000
conv2d_2 0.5000
conv2d_3 0.5000
max_pooling2d_1 0.5000
flatten 0.5000
dense 0.5000
dense_1 0.5000
dense_2 (faulty) 0.5000
Results and Discussion: Table 18 shows that out of 20 training runs, 9 runs failed (i.e., loss did not decrease), while 11 runs passed (i.e., loss decreased). Our findings (see Table 19) show that Tarantula assigns the same suspiciousness score to all layers in the model. Since every layer is activated during the forward and backward passes, there is no variability in their passing/failing test coverage. We also incorporated neuron-level coverage for more granular analysis and visualized the results in the color-coded heat map (see Fig.7). In this plot, each row represents a layer, and each column shows the top ten neurons ranked by descending suspiciousness (from 0.0 to 1.0). We chose the top ten neurons for a concise visualization because our conv2d layers contain 64–128 filters, while the dense layers have up to 256 neurons each (see Fig.6), making it impractical to display every neuron in a single figure.
From Fig.7, we see that specific layers, such as conv2d_2, conv2d_3, and conv2d_4, have suspicious neurons, suggesting their strong association with failing runs. In contrast, conv2d_1 and faulty_output (i.e., the last layer where the bug exists) exhibit high but varying levels of suspiciousness. It should be noted that the final_output layer contains the bug. Although neuron-level coverage reveals these differences, suggesting that not all neurons in a layer behave the same, the overall patterns indicate uniformly high or near-universal activations. Since traditional spectrum-based bug localization techniques rely on such scores to pinpoint the buggy location, the reported trends make bug localization difficult.
Our preliminary study thus confirms the limitation of spectrum-based bug localization techniques, such as Tarantula, in the context of deep learning systems. The results emphasize that while neuron-level coverage offers more detailed insights than layer-level coverage alone, both coverage metrics struggle to isolate the bug in the model.
4.1.4 Preliminary study on dynamic techniques for bug localization
Deep learning systems differ from traditional software systems due to their reliance on high-dimensional data, hardware accelerators, and complex model architectures [36]. Bugs in deep learning systems can arise from diverse sources, including model configurations, training procedures, and runtime dependencies, making their localization particularly challenging. While static analysis techniques have been widely applied to software bug localization, their effectiveness in deep learning systems is often limited, as many issues only manifest during execution [18]. To address this gap, dynamic fault localization techniques leverage runtime analysis to detect faults by monitoring execution traces and identifying abnormal behaviors [18]. Thus, in this study, we systematically evaluate three state-of-the-art dynamic techniques – DeepLocalize [18], DeepDiagnosis [20], and DeepFD [21] – to assess their effectiveness in localizing bugs in deep learning systems. While these techniques leverage runtime analysis to detect and locate bugs, their practical applicability is constrained by several factors, as follows.
Computational and Practical Challenges: Dynamic-based approaches like DeepLocalize [18], DeepDiagnosis [20], and DeepFD [21] rely on runtime data collection, which is computationally expensive and often impractical for large-scale deep learning programs. Moreover, many bug reports in the Denchmark [32] dataset are linked to large, industry-scale deep learning programs that often miss critical information (e.g., dataset details required to run the model) or include numerous external dependencies [37]. These factors make dynamic analysis both time-consuming and, in many cases, impractical.
Nature of the Techniques: Dynamic approaches differ fundamentally from the static, query-based techniques that we employ. These tools (e.g., DeepLocalize [18], DeepDiagnosis [20], DeepFD [21]) first detect and diagnose faults during program execution, identifying issues at a fine-grained level (such as bugs in specific layers or components like loss function). In contrast, traditional static methods use textual and structural information in bug reports to locate files, methods, or lines where bugs exist [30, 17, 11, 12, 64]. Our study focuses on file-level bug localization, which aligns more with developers’ practical steps when addressing a bug, first locating the relevant file in large codebases and then examining its details.
Scope of Bugs Addressed: Another important consideration is that DeepLocalize [18], DeepDiagnosis [20], and DeepFD [21] are specifically designed to address training faults. For example, DeepLocalize [18] targets training-phase numerical faults by analyzing dynamic traces. DeepFD [21] uses machine learning classifiers (e.g., KNN, Decision Tree, and Random Forest) to predict and localize specific training faults (e.g., issues with loss functions, epochs, optimizers, activation functions, or learning rates). DeepDiagnosis [20] focuses on CNN architectures by mapping runtime error symptoms to root causes. Since our dataset, which reflects real-world scenarios, contains a diverse mix of deep learning bugs beyond just training faults, applying these techniques across the entire dataset is neither practical nor appropriate.
Framework and Architecture Limitation: All these techniques are specifically designed to detect and address faults within the TensorFlow/Keras framework only [18, 20, 21]. Moreover, neither DeepLocalize [18] nor DeepDiagnosis [20] supports RNN architectures.
Evaluation of Dynamic Approaches: To ensure a fair and comprehensive evaluation, we followed a structured approach:
  1. 1.
    First, we selected 350 bug reports from our dataset (i.e., Denchmark) using stratified sampling, ensuring that they exclusively use TensorFlow/Keras framework. This subset was chosen to achieve a 95% confidence level with a 5% margin of error, ensuring it is statistically representative of our deep learning system (DLSW) dataset.
  2. 2.
    Second, to better understand the performance of dynamic approaches on the training faults they are designed to handle, we manually identified training bugs within our stratified sample. Among the 350 bug reports, we identified 123 unique DL models. This is because multiple bug reports often correspond to the same DL model in the Denchmark dataset for DLSW. We also excluded the RNN-based DL models since DeepDiagnosis and DeepLocalize do not support them, which left us with 101 DL models.
  3. 3.
    Finally, two independent annotators (the authors) conducted a two-phase labeling process. In the first phase, each DL buggy model was categorized as either a training bug or another type of bug. In the second phase, the training bugs were further classified by their root causes, following the definitions provided by dynamic-based approaches [31]. Before starting the complete annotation, the annotators jointly labeled a sample of 10 buggy DL models to ensure a shared understanding of the classification criteria. The inter-annotator reliability, measured by Cohen’s kappa coefficient, was 0.88, indicating a strong level of agreement. Any disagreements were resolved through a systematic consensus-building process that involved re-examining the cases using clear taxonomy definitions and additional resources such as literature.
This process resulted in 21 unique DL models containing training bugs with well-defined root causes. We then evaluated [18], DeepDiagnosis [20], and DeepFD [21] on these models.
Results and Discussion:
Table 20: Performance of Dynamic Approaches for Locating Training Bugs from DLSW
Technique Accuracy Precision Recall F1-Score
DeepLocalize 0.23 0.25 0.20 0.22
DeepFD 0.28 0.30 0.25 0.27
DeepDiagnosis 0.33 0.35 0.30 0.32
The performance of dynamic approaches across the 21 models reflects the challenges posed by the heterogeneous nature of DL bugs. While these dynamic approaches demonstrate strengths in localizing specific training bugs (e.g., numerical anomalies, learning rate issues), their limitations underscore the multifaceted nature of DL bugs and the inadequacy of existing state-of-the-art techniques for bug localization. Table 20 shows the performance of these techniques against each of the 21 bugs. Based on our evaluation, we found several limitations of these techniques, as follows.
1. Limited Scope: All three dynamic techniques demonstrate significant limitations, even when applied to a small subset of training bugs, due to their narrow scope (see Table 21). DeepLocalize, for instance, focuses on identifying numerical anomalies like NaN or Inf values during bug tracing. While it successfully located a faulty layer caused by an incorrect activation function by detecting NaN values in the loss (Bug 17), it failed to detect bugs that did not involve numerical abnormalities (e.g., Bug 1, 2, 3). Similarly, DeepDiagnosis relies on predefined heuristics to map symptoms to root causes, which restricts its ability to identify issues like an incorrect dropout rate (Bug 21). DeepFD, which uses classifiers trained on runtime features, also struggled with bugs outside its predefined fault categories, such as regularization issues (Bug 10). Our analysis above reveals a troubling trend: even for the specific category of bugs for which they were targeted (e.g., training bugs), these state-of-the-art techniques often fail to detect issues. This suggests that they may not be sufficient for effective bug localization in real-world deep-learning projects, where the bugs tend to be more complex and varied, even within the same category.
2. Inability to Handle Extrinsic Bugs: A key finding from RQ3 is that 40% of DL bugs are extrinsic, meaning they arise from external dependencies such as GPU issues, third-party libraries, or training data. We found that these dynamic bug localization techniques, which rely on analyzing runtime information, are inherently unable to detect such bugs. For example, in our evaluation (see Table 21), the three methods failed to identify bugs caused by extrinsic factors like incorrect or mislabeled training data (Bug 4, 16). These issues originate from issues in the data pipeline rather than the model’s execution, making them invisible to runtime analysis. This limitation highlights a critical gap — extrinsic bugs require approaches that go beyond code execution or runtime monitoring.
3. Challenges with Complex System Interactions: A major insight from our evaluation of training bugs (21 buggy models) is that both static and dynamic methods struggle to address the intricate, often indirect interactions inherent in deep learning systems. For example, state-of-the-art static techniques like bjXnet [30] focus on code and textual bug reports, often overlooking issues such as unchanged model weights that only surface during execution (e.g., Bug 9). Meanwhile, dynamic tools like DeepLocalize, DeepDiagnosis, and DeepFD detect runtime anomalies (e.g., NaN values in Loss) but often fail to link these symptoms to their root causes, such as improper dropout rates causing NaN values in the loss (Table 21). Though dynamic approaches identified some training faults, their overall accuracy remained below 33% (Table 20), highlighting their inability to address deeper systemic interactions.
Our findings demonstrate that deep learning software presents unique challenges that cannot be fully resolved by static or dynamic analysis alone. The limitations observed in both approaches suggest that bugs often arise from complex interactions between code, runtime behaviour, and system-level dependencies, which current tools are not equipped to handle comprehensively. These results align with our broader evaluation, emphasizing the need for more integrated techniques to localize bugs in DLSW effectively.
Table 21: Performance Comparison of Dynamic Approaches for Training Bug Localization
No. Symptom Root Cause (Bug) DL DF DD
1 Loss Not Decreasing Incorrect Loss Function $\times$ $\checkmark$ $\checkmark$
2 Loss Not Decreasing Improper Activation Function $\times$ $\times$ $\checkmark$
3 Loss Increasing After Epochs Incorrect Optimizer $\times$ $\checkmark$ $\times$
4 Model Overfitting Incorrect Training/Validation Data $\times$ $\times$ $\times$
5 Static Validation Loss Incorrect Dropout Usage $\times$ $\times$ $\times$
6 Vanishing Loss & Overfitting Small Training Data $\times$ $\times$ $\times$
7 Overfitting Incorrect Regularizer & Small Data $\times$ $\times$ $\times$
8 Loss Oscillating & Overfitting Incorrect Dropout Rate $\times$ $\times$ $\times$
9 Unchanged Weights & Static Loss Incorrect Learning Rate $\times$ $\checkmark$ $\checkmark$
10 Poor Generalization Incorrect Normalization & No Regularizer $\times$ $\times$ $\times$
11 NaN Values in Loss Incorrect Weight Initialization $\checkmark$ $\times$ $\checkmark$
12 Poor Performance (Low Accuracy) Incorrect Epoch Number $\times$ $\checkmark$ $\times$
13 Train/Val Loss Divergence Incorrect Batch Normalization $\checkmark$ $\times$ $\times$
14 Neurons Outputting Zero Incorrect Weight Initialization $\checkmark$ $\times$ $\checkmark$
15 Training Slowdown Incorrect Dropout Rate $\times$ $\times$ $\times$
16 Overfitting Incorrect Regularizer/Augmentation $\times$ $\times$ $\times$
17 NaN Values in Loss Incorrect Activation Function $\checkmark$ $\checkmark$ $\checkmark$
18 NaN Values in Loss Zero Sample Weights $\checkmark$ $\times$ $\times$
19 Loss Plateaus Improper Bias Initialization $\times$ $\times$ $\times$
20 Performance Degradation Incorrect Learning Rate $\times$ $\checkmark$ $\checkmark$
21 Poor Generalization Incorrect Dropout Rate $\times$ $\times$ $\times$
Accuracy (%) 23.8 28.6 33.3
DL: DeepLocalize, DF: DeepFD, DD: DeepDiagnosis. $\checkmark$: Successfully identified, $\times$: Failed to identify.
@@ 21 Poor Generalization Incorrect Dropout Rate $\times$ $\times$ $\times$
4.2 Answering RQ2: How do different types of bugs in deep learning systems impact bug localization?
In this research question, we investigate the characteristics and localization challenges of different types of bugs in deep learning systems through manual analysis. To classify different types of bugs in deep learning systems, we adopted the following steps:
@@ 4.2 Answering RQ2: How do different types of bugs in deep learning systems impact bug localization?
  • We conducted our manual analysis on a sample of 385 bug reports, which have a 95% confidence level and a 5% margin of error. This allows our findings to be generalizable to the original population of deep learning system bugs.
  • We classified deep learning-related bugs into five categories using the taxonomy of Humbatova et al. [31]. This taxonomy provides explicit definitions for various DL bug types, including Model, Training, Tensor, API, and GPU issues. For instance, bugs involving neural network architecture or weight initialization were classified as ’Model’ bugs, while issues with data preprocessing or augmentation during model training were categorized as ’Training’ bugs. On the other hand, file handling errors or user interface glitches were categorized as Non-deep learning-related (NDL) bugs.
  • Two independent annotators (authors of this study) conducted a two-phase annotation. In the first phase, each bug was labeled as DL or NDL. In the second phase, DL bugs were further categorized into the five sub-types mentioned above. Prior to annotation, the annotators jointly labeled a sample of 20 bug reports to assess their understanding of the classification criteria.
  • The annotators cross-checked their annotations/classifications with project documentation and bug reports to prevent overlaps or misclassifications. Project documentation provided context and clarified the intended system functionality, helping to classify bugs accurately and avoid misclassification.
  • We measured inter-annotator reliability using Cohen’s kappa coefficient [83], resulting in a score of 0.80. This substantial agreement indicates a high level of consistency in our classification process. We resolved disagreements through a systematic consensus-building process involving joint re-examination of disputed cases, guided by taxonomy definitions and cross-referenced with additional documentation to ensure accuracy and consistency.
  • We documented this manual analysis using an Excel sheet, with a total of $\approx$55 hours spent by each author, which is provided in our replication package [66].
@@
Figure 9: Performance of existing bug localization techniques (BugLocator, BLUiR, BLIA, DNNLOC, bjXnet) for each type of bug in deep learning systems
Localization of bugs in deep learning systems: To gain a deeper understanding of the challenges in localizing deep learning bugs, we analyzed the performance of our baseline techniques across different categories of bugs in DL systems. We employed stratified random sampling [84] to select a balanced number of bug reports (100 bugs) from each category for performance comparison, ensuring fair representation and reducing selection bias. We repeated the process three times with different samples to ensure robustness. The results were averaged across these evaluations, providing reliable insights into the efficacy of the bug localization methods. In addition to the quantitative results, we conducted an in-depth analysis of each type of DL bug to understand their inherent challenges and how they impact the overall performance of the bug localization techniques. This detailed analysis, with examples, is provided below.
Table 22: Experimental result of existing bug localization techniques (BugLocator, BLUiR, BLIA, DNNLOC, bjXnet) for each category of bugs in deep learning systems
Method Model Training Tensor API GPU MAP BugLocator 0.368 0.312 0.235 0.446 0.183 BLUiR 0.293 0.403 0.601 0.258 0.266 BLIA 0.357 0.345 0.448 0.395 0.290 DNNLOC 0.593 0.411 0.619 0.493 0.296 bjXnet 0.650 0.420 0.650 0.460 0.310 MRR BugLocator 0.532 0.386 0.358 0.378 0.223 BLUiR 0.387 0.427 0.621 0.424 0.311 BLIA 0.472 0.419 0.553 0.447 0.327 DNNLOC 0.585 0.467 0.682 0.548 0.336 bjXnet 0.720 0.480 0.680 0.510 0.360 MAP= Mean Average Precision MRR=Mean Reciprocal Rank
Model bugs: Model bugs present unique challenges in bug localization for deep learning systems. Our analysis in Table 22 reveals that bjXnet demonstrates superior performance in localizing model bugs. This can be attributed to its attention mechanisms, which enable focused examination of key subcomponents. Model bugs typically relate to a model’s type, properties, and layers. We observed significant vocabulary overlap between bug reports describing model-related issues and the implementation of deep learning models. bjXnet’s attention-based approach allows it to prioritize critical code sections, such as specific methods or functions directly implicated in the bug report. DNNLOC also performs well, followed closely by bjXnet, by employing deep neural networks to capture complex patterns and utilizing rVSM to assess textual relevance between bug reports and source code. In contrast, BLUiR shows the poorest performance in locating model bugs.
@@
Table 23 [40] illustrates an example of a model bug from the CASL.ai project, with the corresponding code snippet shown in Fig. 12 [40] (Appendix A). The bug occurs in the bpe method of the GPT2Tokenizer, causing faulty tokenization that impacts the GPT2 language model’s functionality.
We noticed that IR-based techniques struggle with locating model bugs. For instance, BugLocator incorrectly identifies ’SentencePieceTokenizer.py’ (Fig. 13, Appendix A) as the Top@1 result, which is likely to stem from lexical overlap between the bug report and source code. Four key terms from the bug report (Table 23) – GPT2Tokenizer, recover, seq2seq, and model – appear in the incorrect file, potentially leading to this incorrect localization. Notably, BugLocator retrieves the actual ground truth file at the 8th position.
On the other hand, BLIA ranks the ground truth relatively low at the 17th position, potentially because meta components, such as stack trace, version control history, or commit history, are less effective in localizing model bugs. Model bugs typically involve deep architectural issues within the neural network, which are not easily traced through surface-level meta-information, making BLIA less suitable for these types of bugs compared to traditional software bugs.
@@
BLUiR performs even worse, ranking the ground truth code at the 131st position. This suggests that the similarity between bug reports and source code elements may be insufficient for accurately locating model bugs. In the example from Table 23, the bug manifests when tokenizing the sentence ’BART is a seq2seq model.’ BLUiR mistakenly retrieves the source code file containing the Seq2Seq class as the Top@1 result. This error likely occurs because several important keywords from the bug report align with similar code elements (e.g., ’seq2seq’, ’encode’, and ’model’ matching the seq2seq class). Such misalignment results from BLUiR’s strong emphasis on structural elements (e.g., class, method, variable) and insufficient consideration of the bug report’s semantic aspects.
In contrast, both bjXnet and DNNLOC successfully retrieve the buggy file at Top@1. Unlike IR-based methods, DNNLOC’s hybrid approach leveraging neural networks can capture deeper semantic links beyond lexical similarities. For instance, despite textual mismatches, it can better connect terms like ’seq2seq’ and ’tokenizer’ in the bug report to their corresponding code-level implementations in GPT2Tokenizer. DNNLOC’s ability to leverage non-linear relationships likely contributes to its superior performance in localizing model bugs in deep learning systems. bjXnet’s success can be primarily attributed to its attention mechanism, which enables the targeting of critical subcomponents within the code, such as the bpe method in the GPT2Tokenizer. By focusing on the most relevant methods and variables implicated in the bug report, bjXnet achieves higher precision, especially in modular deep-learning components like tokenization and layer definitions, where pinpoint accuracy is crucial for fault identification.
Training bugs: All five approaches perform poorly in localizing training bugs, as shown in Table 22. bjXnet, DNNLOC, and BLUiR perform comparatively better than the other techniques, though their overall MAP values remain low (40.30% – 42.00%). Table 24 [41] illustrates a training bug from the fast.ai project, where the interaction between Gradient Accumulation and the MixedPrecision Callback leads to artificially high training loss values.